two positive point charges q are placed on the y-axis at y=a and y=-a. a negative point charge "-q" is located at some point in the +-axis.

1
Expert's answer
2022-04-11T12:02:11-0400

"F=2F_x=2\frac{kQq}{x^2+a^2}\cdot \frac x{\sqrt{x^2+a^2}}=\frac{2kQqx}{\sqrt{(x^2+a^2)^3}}."